1. (9.19) A flywheel turns at 480 rpm. Compute the angular speed at any point on the wheel and
the tangential speed if it has a radius of 30 cm. Note: tangential speed can be considered as the
linear speed of the wheel.
3. Compute the rotational Kinetic energy of a 25 kg wheel rotating at 6 rev/s if the radius of
gyration of the wheel is 22 cm.
Note: the radius of gyration is the point where the axis of rotation passes through. This axis is a
result of the parallel axis theorem. In this problem, since it is given, you can already compute for
what is being asked for.
4. To crack a certain nut in a nutcracker, forces with magnitudes of at least 40 N must act on its
shell from both sides. For the nutcracker of Fig. 12-34, with distances
1
Expert's answer
2021-07-15T12:42:50-0400

"w=2\pi n=2\pi\times \frac{480}{60}=16\pi"

Anguler velocity


"V=wr"


"V=16\pi\times0.30=15.072m/sec"


"I=\frac{1}{2}mr^2=\frac{1}{2}\times25\times0.22^2=0.605kgm^2\\w=2\pi\times6=12\pi rad/sec"

"K=\frac{1}{2}Iw^2=\frac{1}{2}\times0.605\times(12\pi)^2=430J"


L=12cm

d=2.6cm

To crack a certain nut in a nutcracker, forces with magnitudes of at least 40 N

"F'=F.\frac{L}{d}=40\times\frac{2.6}{12}=9N"